A boxcar of length 9.1 $m$ and height 2.4 $m$ is at rest on frictionless rails. Inside the boxcar (whose mass when empty is 3900 $kg$) a tank containing 1400 $kg$ of water is located at the left end. The tank is 0.8 $m$ long and 2.4 $m$ tall. At some point the walls of the tank start to leak, and the water fills the floor of the boxcar uniformly. Assume that all the water stays in the boxcar. After all the water has leaked out what will be the final velocity of the boxcar? answer: 0.0 m/s (Take movement to the right as positive. Assume that the mass of the boxcar is evenly distributed.) What is the displacement of the boxcar 7 $s$ after the water has settled in the bottom. (Take positive displacement as being to the right.)
Answer #1

Finding the COM of the system, we get

Xcom = (3900 x 9.1/2 + 1400 x 0.8/2)/(3900 + 1400) = 3.45 m

from conservation of linaer momentum the COM should stay at the same location for an external distant observer, so

D = 9.1/2 - 3.45 = 1.1 m,

Hence, D = 1.1 m, to the left of its intial position.